OpenStudy (anonymous):

solve by substitution; y=8x+23 y=-4x-13 Please/Thanks:)

OpenStudy (anonymous):

-4x -13 = 8x + 23

OpenStudy (anonymous):

can you solve for x above ?

OpenStudy (anonymous):

EASY bro!!! we simply realize that if y=*x+23 then we can say that 8x+23= -4x -13. if we solve this simple equation, we see that x= -4. from there simply put this into y= 8x +23 and see that y= -13

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

:)